Circular No. 6106

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

X-RAY NOVA IN SCORPIUS S. N. Zhang, B. A. Harmon, W. S. Paciesas, C. A. Wilson, and G. J.

Fishman report for the Compton Gamma Ray Observatory BATSE Team: “The intensity of the current outburst from this object (cf. IAUC 6101) has gradually increased from 710 mCrab on Nov. 7 to a peak of ∼ 830 mCrab between Nov. 10 and 11. It has since declined to 660 mCrab as of Nov.

15. The approximate error on these intensities is 5 percent. The photon spectrum is well fit with a power law of index −3.1 ± 0.1 during this period. Immediate follow-up radio observations of this source are strongly encouraged, because a giant radio outburst was reported (IAUC 6055) during the decline of the hard x-ray flux in August (IAUC 6046 and 6056). Relativistic jets were formed shortly after that (IAUC 6060).”

PKS 0422+004 M. Fiorucci and G. Tosti, Astronomical Observatory and Department

of Physics, University of Perugia, report that their photoelectric observations show that this BL-Lac object is in a high-luminosity stage with superimposed rapid variations having temporal scales of a week or less. CCD V magnitudes: Oct. 15.10 UT, 15.20; Nov. 5.08, 14.57; 11.98, 15.26; 17.98, 13.85; 18.02, 13.79.

V1333 AQUILAE M. Garcia, P. Callanan, J. McClintock, and P. Zhao, Smithsonian As-

trophysical Observatory report: “V1333 Aql, the optical counterpart of the x-ray binary Aquila X-1, is once again in its low state. CCD images obtained between 1994 Apr. 9 and July 31 show that the V magnitude ranged from 17.4 to 18.0. However, our most recent observations on Nov. 5.1 and 14.1 UT found V = 18.96 and 19.30 (± 0.04), respectively.”

PERIODIC COMET BORRELLY (1994l) Some observers have commented on the existence of a significant fan-

shaped anti-tail around Nov. 7–13 UT. Total visual magnitude estimates by C. S. Morris from southern California (L = 0.26-m reflector; B = 20×80 binoculars): Oct. 2.47 UT, 9.0 (L); 3.47, 8.9 (L); 7.47, 8.8 (L); 8.48, 8.8 (L); 9.48, 8.8 (L); 12.44, 9.0 (L); 30.36, 8.0 (B); Nov. 12.50, 7.9 (B).

1994 November 19 Daniel W. E. Green